4D Flat on DX 20180309

17:30 4D image is saturated. Probably communication problem between 4D computer and 4D Power Supply

17: 50 Leroy turn off laser, power supply, computer. Turn on computer, wait, power supply and laser.
  • Still saturated

18:00 Leroy turn off laser, power supply, computer. Turn on computer, wait, power supply and laser.
  • Still saturated

18:10 When Leroy turned off the 4D laser, the intensity went to 0. The implication is that the 4D camera was, by default because a communication problem, set to 100 ms (or something) and thus the image was saturated.
  • Leroy will check if there is a main power switch on the back of the 4D power supply.
  • Works. Leroy has pulled the power cord from the back of the power supply and the laser box. Great!

18:40 Aligned manually with hexapod commands, focus and coma.

18:50 Started 4D server and AdSec client. Tested with a=get_img4d(). Worked

18:55 print, hexalign()
  • applied manually. Looks much better

19:00 print, hexalign(/apply)
  • removed a bit of tip/tilt.
  • maybe astigmatism left and high order

19:40 Default flat applied
  • had ~100 nm rms before

19:44 Flattened with 200 modes
  • 58 nm
  • still see a lot of high order along outside edge

19:50 400 modes
  • 45 nm

19:54 600 modes
  • 39 um rms
  • flat 20180309_195423
  • New flat: Interferometer
    full_600modes_inter.jpg
  • New flat: Wavefront with T/T/Focus/Coma removed
    full_600modes_low_removed.jpg
  • New flat: Applied Current
    full_600modes_current.jpg

20:10 500 modes
  • 52 um rms

20:17 Noise data (HBS On)
  • 201700_noise
  • 201900_noise
  • 202000_noise
  • 202100_noise

20:22 Rest and apply flat 20180309_195423
  • 639 modes
  • 90 um rms (41 um with astig removed)
  • 97 um rms (38 um with astig removed)
  • 92 um rms (41 um with astig removed)

20:37 Apply 600 mode correction and hexalign
  • 40 um rms ( 39 um with astig removed)

20:38 Noise data (HBS On)
  • 203800_noise
  • 203900_noise
  • 204000_noise

20:50 Tracking numbers
tn = ['20180309_201700_noise','20180309_201900_noise','20180309_202000_noise','20180309_202100_noise','20180309_203800_noise','20180309_203900_noise','20180309_204000_noise']

20:50 Done, Elliott will fill Mods and LBTI while Greg sets up LUCI2 for TT LUT measurement

-- %USERSIG{DougMiller - 2018-03-09}%

Comments

 
Topic revision: r3 - 09 Mar 2018, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback